Applications of Differentiation

Consider the function f(x)= cos(x) - (1/2)x.  This function has two critical numbers  A<B in [0,2pi]. Give the following:    

A=

B=

f''(A)=

f''(B)=
Answer #1

f(x)=cosx +(√2 /2)x

f '(x)=-sinx+(√2 /2)

for critical numbers f '(x)=0

-sinx+(√2 /2)=0

sinx=(√2 /2)

x=pi/4 ,3pi/4

A=pi/4

B=3pi/4

f "(x)=-cosx

f "(pi/4)=-cos(pi/4)= -(√2 /2) <0

f "(3pi/4)=-cos(3pi/4)= (√2 /2) >0

f(x) has a localmaximum at A and a local minimum at B


answered by: ANURANJAN SARSAM